2kclo3------2kcl+3o2,5.0 g of kclo3 gave0.03 mol  of o2 hence the percenatge purity of kclo3 is 

Grade:

1 Answers

Sunil Kumar FP
9 years ago

%age purity can be determined as follows.
suppose u hav 'x' gm pure copound and 'y' gm impure substance or inert material, then the purity %age of x is given by
(x/x+y)*100

3 mole of o2 give 2 mole kclo3
mole of kclo3 in 5 g=5/122.5=.04
now .03 mole will give=.02 mole
pure kclo3 in gm=.02*122.5=2.45
therefore the percentage purity is 2.45/5*100=49%